# False Discovery Rate and Localizing Power

### v2

@@ Abstract
False discovery rate (FDR) is commonly used for correction for multiple testing in neuroimaging studies. However, when using two-tailed tests, making directional inferences about the results can lead to vastly inflated error rate, even approaching 100% in some cases. This happens because FDR control the error rate only globally, over all tests, not within subsets, such as among those in only one or another direction. Here we consider and evaluate different strategies for FDR control with two-tailed tests, using both synthetic and real imaging data. Approaches that separate the tests by direction of the hypothesis test, or by the direction of the resulting test statistic, more properly control the directional error rate and preserve FDR benefits, albeit with a doubled risk of errors under complete absence of signal. Strategies that combine tests in both directions, or that use simple two-tailed p-values, can lead to invalid directional conclusions, even if these tests remain globally valid. To enable valid thresholding for directional inference, we suggest that imaging software should allow the possibility that the user sets asymmetrical thresholds for the two sides of the statistical map. While FDR continues to be a valid, powerful procedure for multiple testing correction, care is needed when making directional inferences for two-tailed tests, or more broadly, when making any localized inference.
@@
False discovery rate, two-tailed tests, localizing power, subsetting property
@@ 1 Introduction
In the context of the multiple testing problem, control over the false discovery rate (fdr) provides an alternative to the control over the familywise error rate (fwer): while fwer refers to the chance of even one type i (false positive) error is committed in a set (family) of statistical tests, fdr refers to the proportion of such errors among tests for which the null hypothesis has been rejected, that is, the proportion of false discoveries among all discoveries. These topics have been extensively covered in the literature; for an early review of the multiple testing problem in brain imaging, see Nichols and Hayasaka (2003); for a review in genetics, also broadly applicable to imaging, see Goeman and Solari (2014); for early papers and for an overview of the development of fdr, see Seeger (1968); Sorić (1989); Benjamini and Hochberg (2000); for its introduction to neuroimaging, see Genovese et al. (2002); for a critique, see Chumbley and Friston (2009); for a discussion of various fdr approaches, see Korthauer et al. (2019). Methods for fwer correction in brain imaging are often based on the random field theory (rftWorsley et al., 1996, 2004), permutation tests (Holmes et al., 1996; Nichols and Holmes, 2002; Winkler et al., 2014), Monte Carlo simulations (Poline and Mazoyer, 1993; Forman et al., 1995; Cox, 1996) and occasionally, the method based on Bonferroni (1936).
@@
Two hallmarks of most neuroimaging studies are that they can involve thousands of statistical tests, and that one can reasonably expect effects to be observed in both directions for many of these. Controlling the fdr aims to provide more power than fwer in the setting where many null hypotheses are false. This benefit has seen fdr widely adopted in neuroimaging. However, the behavior of fdr when signal manifests in positive direction for some tests, and negative for others, has not received as much attention, even if it is a pertinent problem in the field. In principle, using fdr with two-tailed tests is straightforward: the two-tailed p-values are obtained, then an fdr-controlling procedure is applied. As long as inference remains focused on all tests, using fdr on two-tailed p-values has nothing special compared to using one-tailed p-values. Using two-tailed tests can become problematic if, after doing the test, the researcher proceeds to draw inferences on each direction: fdr is not controlled within one direction or another, only globally. Previous literature on directional hypotheses pertinent to this problem can be found in Shaffer (1995, 2002); Williams et al. (1999); Benjamini and Yekutieli (2005). The same problem emerges if a researcher attempts to draw inferences within regions of a map of p-values after applying fdr globally across all of them. Both problems are a general manifestation of the fact that fdr, while a powerful procedure, does not have the ability to allow localization within subsets of tests, that is, it lacks localizing power. As one of our reviewers pointed out, “making discoveries in either direction while controlling the fdr is trickier than it may seem”.
@@
In this note we start by briefly revisiting the multiple testing problem in imaging, p-value adjustment, the use of fdr, and discuss interpretation for two-tailed tests in the light of statistical properties of fdr-controlling procedures. We demonstrate that merely applying fdr does not control the error rate on each side of the statistical map when the researcher wishes to make inferences about the direction after using two-tailed tests, and can lead to invalid conclusions when effects are asymmetrically distributed across the two sides. We investigate methods that can be considered to address such issues, with focus on a slight modification of the method proposed by Benjamini and Yekutieli (2005). We also bring attention to the fdr approach proposed by Benjamini et al. (2006), which is more powerful than the one originally proposed by Benjamini and Hochberg (1995), and which has seldom been used in brain imaging. We also offer a method for p-value adjustment using this more powerful procedure. While the paper is presented in a neuroimaging context, the same topics are relevant to other fields in which fdr is widely used.
@@ 2.3 The basic fdr procedure: bh
Control over the fdr can be achieved with the procedure by Benjamini and Hochberg (1995) (bh): let $p_{(1)}\leqslant\ldots\leqslant p_{(V)}$ be the ordered p-values. Find the largest index $i$ for which:
@@
When the tests are independent, the bh procedure controls the fdr at the level $qV_{0}/V$, where $V_{0}$ is the (unknown) number of true null hypotheses. However, complete independence is not strictly necessary: Benjamini and Yekutieli (2001) showed that it is sufficient that tests exhibit a form of dependence that they termed positive regression dependency on subsets (prds). In this form of dependence, knowledge that one p-value where the null is true is small does not make less likely that the other (V1)(V-1) p-values are also small. If this condition is satisfied, the bh procedure still controls the fdr. Analyses where a false positive does not make the other tests less likely to be significant, as typical in brain imaging, satisfy the prds condition.222prds itself is defined in Benjamini and Yekutieli (2001); additional literature on forms of dependency between variables, relevant to prds, can be found in Lehmann (1966); Esary et al. (1967); Sarkar (1969); Kimeldorf and Sampson (1989).
A procedure that controls the fdr more generally, for any dependence between tests (as opposed to only under independence or prds) was proposed by Benjamini and Yekutieli (2001) (by). The procedure is similar to bh, except for a small modification in Equation 1, which becomes $p_{(i)}\leqslant iq/V/c(V)$, where $c(V)=\sum^{V}_{j=1}1/j$. Graphically, this results in a less inclined threshold line, and thus, a smaller cutoff $p_{(i)}$ (more conservative) than with the original bh procedure.
2.4 A more powerful fdr procedure: bky
Unfortunately, bh and by are both conservative in the sense that the expected fdr is below $qV_{0}/V$, which is always less than or equal to the desired level $q$. To mitigate this conservativeness, Benjamini et al. (2006) proposed to estimate $V_{0}$ after a first pass using bh with $q^{\prime}=q/(1+q)$, then repeat with $q^{*}=q^{\prime}V/\hat{V}_{0}$ if at least one hypothesis was rejected in the first pass; this two-pass procedure has been proposed for use in brain imaging (Chen et al., 2009), albeit it remains rarely used. In their original paper, Benjamini et al. (2006) recognised that there is no reason to stop after the second pass; in fact, the procedure can be repeated iteratively until there are no more rejections. Doing so leads to following procedure (bky), which we adopt here: starting from the lowest p-value, keep rejecting hypotheses for the successive $p_{(i)}$ as long as at least one $p_{(j)}$, $j\geqslant i$, satisfies:
$p_{(j)}\leqslant\frac{jq}{V+1-i(1-q)}$ (2)
stopping at the first $i$ for which the inequality is no longer satisfied for any $j\geqslant i$; all subsequent nulls, inclusive, are retained (Benjamini et al., 2006, Definition 7).
@@ 2.5 Adjustment of p-values
Unfortunately, corrected p-values for bh and bky are not guaranteed to be monotonically related to their corresponding uncorrected p-values (that is, corrected p-values not necessarily increase or decrease as their corresponding uncorrected p-values increase or decrease). Graphically, the corrected p-values correspond to the slope of the line that connects each point to the center of the Cartesian grid, and that slope can be higher than $q$ even for p-values considered significant. Such lack of monotonicity is problematic because thresholding a map of such corrected p-values at some desired level $q$ will not produce the same result as thresholding the map of original p-values at the critical $p_{(i)}$ identified with the respective fdr-controlling method. One may expect monotonicity when doing statistical tests: it allows one to trade-off between maximizing power while minimizing false positives (Zeisel et al., 2011). Following Yekutieli and Benjamini (1999, Definition 2.4), adjusted p-values, which replace corrected p-values without having this issue, can be computed for the bh procedure as:
@@
With adjusted p-values, significance thresholds no longer depend on the position $i$ for bky: maps of adjusted p-values can be arbitrarily thresholded at any $q$ in the interval between 0 and 1 (Figure 2). It should be noted that ties are very common with adjusted p-values; they pose no problem for interpretation. Interestingly, using Equation 6 for bky, as opposed to a formulation equivalent to Equation 5, solves the issue with the paradoxical results observed by Reiss et al. (2012), whereby adjusted p-values would seem smaller than their unadjusted counterparts.
@@ 2.6 Localizing power
Methods for fwer and fdr control differ not only on the measurement of error that is controlled, but on at least another aspect: if a procedure controls the fdr at some level for a set of hypotheses, this control does not necessarily extend to every subset of hypotheses, and the actual fdr can in fact be much higher (or even much lower) for any subset; a procedure that controls the fwer for a set of hypotheses, however, also controls the fwer for every subset. This feature, that fwer procedures have, but fdr procedures have not, was described by Finner and Roters (2001), and called subsetting property by Goeman and Solari (2014). In the context of imaging, we find the term localizing power (or the lack of it) preferable as it immediately conveys the meaning that further localization within parts of an image may or may not be possible, depending on the procedures used for inference (Friston et al., 1996; Nichols and Hayasaka, 2003; Noble et al., 2022).
@@
An investigation into the association between cortical thickness (measured at every vertex of a surface representation of the brain) and a behavioral measure involves a sample of hundreds of individuals. Acknowledging that effects could be manifest in either direction, and pursuing statistical rigor, the researcher employs two-tailed p-values, with adjustments for multiple testing using an fdr. The map of test statistics is then masked at the significant level $q=0.01$, thus retaining for visualization only the significant vertices. The analysis uncovers large areas in the frontal and parietal lobes with a positive relationship to the behavioral measure, as hypothesized. However, it also reveals a negative association spanning broad portions of the fusiform gyrus, as well as vertices showing positive or negative associations scattered across the cortex. The researcher conducting the analysis writes a report detailing the novel findings in the negative direction, emphasizing the stringent methodologies that incorporated two-tailed tests and a strict fdr significance threshold, well below the conventional level $q=0.05$, and concludes (erroneously) that only about 1% of the negative associations are expected to be false positives.
@@
In a study of a compound for reducing the severity of relapses in multiple sclerosis (ms), white matter hyperintensities (wmh) are analyzed voxelwise. Subjects are randomly allocated into active compound and placebo groups. Based on prior knowledge and on random population sampling, normality is assumed and a two-sided group comparison is conducted using a two-sample $t$-test. This produces two-tailed p-values calculated from the Student’s $t$-distribution. Control for multiple testing across voxels is achieved using fdr at $q=0.05$. After determining the fdr threshold, the inverse cumulative distribution function of the Student’s $t$-distribution is used to derive the critical threshold $t_{\textsc{fdr}}$. Given the two-tailed nature of the test, symmetrical thresholds $t_{\textsc{fdr}}$ and $t_{\textsc{fdr}}$ are set in an image visualization tool. The results show a notable reduction in the periventricular wmh load in the group that used the active compound, as expected (and hoped), as well as mixed results in various juxtacortical regions, albeit with mostly increases in wmh load in these areas. In the report, the investigator concludes (erroneously) that the drug may amplify the wmh load in juxtacortical areas given that, as fdr was used, no more than 5% of the observed changes are false positives.
@@
These issues are a consequence of the lack of the subsetting property. Localization errors can be avoided by refraining from making statements or taking actions that could refer to the amount of errors in specific regions when fdr is controlled only globally, or by doing separate fdr analyses within each subset (Efron, 2008). The same holds to avoid direction errors. Direction errors can also be avoided, we assert, by partitioning the statistical map into two disjoint sets: one with the positive test statistics, another with the negative test statistics, then proceeding to fdr correction in each subset using their two-tailed p-values. This is expected to work because under the null hypothesis, the partitioning using the sign of the test statistic is independent of the two-tailed p-values; the latter are distributed uniformly between 0 and 1 and, in either side of the map, more extreme test statistics (stronger effects) imply lower p-values. Thus, the requirements for fdr, using either bh or bky, are satisfied. The fdr can then be controlled at the desired level $q$ for the positive and for the negative sides of the map. We verify this in the next section, along with other methods.

### v3

@@ Abstract
False discovery rate (FDR) is commonly used for correction for multiple testing in neuroimaging studies. However, when using two-tailed tests, making directional inferences about the results can lead to a vastly inflated error rate, even approaching 100% in some cases. This happens because FDR controls the error rate only globally, over all tests, not within subsets, such as among those in only one or another direction. Here we consider and evaluate different strategies for FDR control in such cases, using both synthetic and real imaging data. Approaches that separate the tests by direction of the hypothesis test, or by the direction of the resulting test statistic, more properly control the directional error rate and preserve FDR benefits, albeit with a doubled risk of errors under complete absence of signal. Strategies that combine tests in both directions, or that use simple two-tailed p-values, can lead to invalid directional conclusions, even if these tests remain globally valid. A solution to this problem is through the use of selective inference, whereby positive and negative tails are treated as sets (families), which are screened locally, then subjected to FDR at a modified level that controls average FDR over those that survive the initial screening. Moreover, the BKY procedure can be used in place of the well-known Benjamini-Hochberg, yielding additional power. These methods are easy to implement. Finally, to enable valid thresholding for directional inference, we suggest that imaging software should allow the user to set asymmetrical thresholds for the two sides of the statistical map. While FDR continues to be a valid, powerful procedure for multiple testing correction, care is needed when making directional inferences for two-tailed tests, or more broadly, when making any localized inference.
@@
False discovery rate, two-tailed tests, localizing power, subsetting property, selective inference, multiple families
@@ 1 Introduction
In the context of the multiple testing problem, control over the false discovery rate (fdr) provides an alternative to the control over the familywise error rate (fwer): while fwer refers to the chance of even one type i (false positive) error is committed in a set (family) of statistical tests, fdr refers to the proportion of such errors among tests for which the null hypothesis has been rejected, that is, the proportion of false discoveries among all discoveries. These topics have been extensively covered in the literature; for an early review of the multiple testing problem in brain imaging, see Nichols2003; for a review in genetics, also broadly applicable to imaging, see Goeman2014; for early papers and for an overview of the development of fdr, see Seeger1968; Soric1989; Benjamini2000; for its introduction to neuroimaging, see Genovese2002; for a critique, see Chumbley2009; for a discussion of various fdr approaches, see Korthauer2019. Methods for fwer correction in brain imaging are often based on the random field theory (rftWorsley1996; Worsley2004), permutation tests (Holmes1996; Nichols2002; Winkler2014), Monte Carlo simulations (Poline1993; Forman1995; Cox1996) and occasionally, the method based on Bonferroni1936.
@@
Two hallmarks of most neuroimaging studies are that they can involve thousands of statistical tests, and that one can reasonably expect effects to be observed in both directions for many of these. Controlling the fdr aims to provide more power than fwer in the setting where many null hypotheses are false. This benefit has seen fdr widely adopted in neuroimaging. However, the behavior of fdr when signal manifests in positive direction for some tests, and negative for others, has not received as much attention, even if it is a pertinent problem in the field. In principle, using fdr with two-tailed tests is straightforward: the two-tailed p-values are obtained, then an fdr-controlling procedure is applied. As long as inference remains focused on all tests, using fdr on two-tailed p-values has nothing special compared to using one-tailed p-values. Using two-tailed tests can become problematic if, after doing the test, the researcher proceeds to draw inferences on each direction: fdr is not controlled within one direction or another, only globally. Previous literature on directional hypotheses relevant to this problem can be found in Shaffer1995; Shaffer2002; Williams1999; Benjamini2005. The same problem emerges if a researcher attempts to draw inferences within regions of a map of p-values after applying fdr globally across all of them. Both problems are a general manifestation of the fact that fdr, while a powerful procedure, does not have the ability to allow localization within subsets of tests, that is, it lacks localizing power. As one of our reviewers pointed out, “making discoveries in either direction while controlling the fdr is trickier than it may seem”.
@@
In this note we start by briefly revisiting the multiple testing problem in imaging, p-value adjustment, the use of fdr, and discuss interpretation for two-tailed tests in the light of statistical properties of fdr-controlling procedures. We demonstrate that merely applying fdr does not control the error rate on each side of the statistical map when the researcher wishes to make inferences about the direction after using two-tailed tests, and can lead to invalid conclusions when effects are asymmetrically distributed across the two sides. We investigate methods that can be considered to address such issues, emphasizing the method proposed by Benjamini2014 for selective inference. We also bring attention to the fdr approach proposed by Benjamini2006, which is more powerful than the one originally proposed by Benjamini1995, and which has seldom been used in brain imaging. We also offer a method for p-value adjustment using this more powerful procedure. While the paper is presented in a neuroimaging context, the same topics are relevant to other fields in which fdr is widely used.
@@ 2.3 The basic fdr procedure: bh
Control over the fdr can be achieved with the procedure by Benjamini1995 (bh): let $p_{(1)}\leqslant\ldots\leqslant p_{(V)}$ be the ordered p-values. Find the largest index $i$ for which:
@@
When the tests are independent, the bh procedure controls the fdr at the level $qV_{0}/V$, where $V_{0}$ is the (unknown) number of true null hypotheses. However, complete independence is not strictly necessary: Benjamini2001 showed that it is sufficient that tests exhibit a form of dependence that they termed positive regression dependency on subsets (prds). In this form of dependence, knowledge that one p-value where the null is true is small does not make less likely that the other (V1)(V-1) p-values are also small; additional literature on forms of dependency relevant to prds can be found in Lehmann1966; Esary1967; Sarkar1969; Kimeldorf1989. Analyses where a false positive does not make the other tests less likely to be found significant, as typical in brain imaging, satisfy the prds condition. If prds is satisfied, the bh procedure still controls the fdr.
A procedure that controls the fdr more generally, for any dependence between tests (as opposed to only under independence or prds) was proposed by Benjamini2001 (by). The procedure is similar to bh, except for a small modification in Equation 1, which becomes $p_{(i)}\leqslant iq/V/c(V)$, where $c(V)=\sum^{V}_{j=1}1/j$. Graphically, this results in a less inclined threshold line, and thus, a smaller cutoff $p_{(i)}$ (more conservative) than with the original bh procedure.
2.4 A more powerful fdr procedure: bky
Unfortunately, bh and by are both conservative in the sense that the expected fdr is below $qV_{0}/V$, which is always less than or equal to the desired level $q$. To mitigate this conservativeness, Benjamini, Krieger, and Yekutieli proposed to estimate $V_{0}$ after a first pass using bh with $q^{\prime}=q/(1+q)$, then repeat with $q^{*}=q^{\prime}V/\hat{V}_{0}$ if at least one hypothesis was rejected in the first pass (Benjamini2006, or simply bky). This two-pass procedure has been proposed for use in brain imaging (Chen2009), albeit it remains rarely used. In their original paper, Benjamini2006 recognised that there is no reason to stop after the second pass; in fact, the procedure can be repeated iteratively until there are no more rejections. Doing so leads to following procedure (bky), which we adopt here: starting from the lowest p-value, keep rejecting hypotheses for the successive $p_{(i)}$ as long as at least one $p_{(j)}$, $j\geqslant i$, satisfies:
$p_{(j)}\leqslant\frac{jq}{V+1-i(1-q)}$ (2)
stopping at the first $i$ for which the inequality is no longer satisfied for any $j\geqslant i$; all subsequent nulls, inclusive, are retained (Benjamini2006, Definition 7).
@@ 2.5 Adjustment of p-values
Unfortunately, corrected p-values for bh and bky are not guaranteed to be monotonically related to their corresponding uncorrected p-values (that is, corrected p-values not necessarily increase or decrease as their corresponding uncorrected p-values increase or decrease). Graphically, the corrected p-values correspond to the slope of the line that connects each point to the center of the Cartesian grid, and that slope can be higher than $q$ even for p-values considered significant. Such lack of monotonicity is problematic because thresholding a map of such corrected p-values at some desired level $q$ will not produce the same result as thresholding the map of original p-values at the critical $p_{(i)}$ identified with the respective fdr-controlling method. One may expect monotonicity when doing statistical tests: it allows one to trade-off between maximizing power while minimizing false positives (Zeisel2011). Following Yekutieli1999, adjusted p-values, which replace corrected p-values without having this issue, can be computed for the bh procedure as:
@@
With adjusted p-values, significance thresholds no longer depend on the position $i$ for bky: maps of adjusted p-values can be arbitrarily thresholded at any $q$ in the interval between 0 and 1 (Figure 2). It should be noted that ties are very common with adjusted p-values; they pose no problem for interpretation. Interestingly, using Equation 6 for bky, as opposed to a formulation equivalent to Equation 5, solves the issue with the paradoxical results observed by Reiss2012, whereby adjusted p-values would seem smaller than their unadjusted counterparts.
@@ 2.6 Localizing power
Methods for fwer and fdr control differ not only on the measurement of error that they target, but on at least another aspect: if a procedure controls the fdr at some level for a set of hypotheses, this control does not necessarily extend to every subset of hypotheses, and the actual fdr can in fact be much higher (or even much lower) for any subset; a procedure that controls the fwer for a set of hypotheses, however, also controls the fwer for every subset. This feature, that fwer procedures have, but fdr procedures have not, was described by Finner2001, and called subsetting property by Goeman2014. In the context of imaging, we find the term localizing power (or the lack of it) preferable as it immediately conveys the meaning that further localization within parts of an image may or may not be possible, depending on the procedures used for inference (Friston1996; Nichols2003; Noble2022).
@@
An investigation into the association between cortical thickness (measured at every vertex of a surface representation of the brain) and a laboratory measure involves a sample of hundreds of individuals. Acknowledging that effects could be manifest in either direction, and pursuing statistical rigor, the researcher employs two-tailed p-values, with adjustments for multiple testing using an fdr. The map of test statistics is then masked at the significant level $q=0.01$, thus retaining for visualization only the significant vertices. The analysis uncovers large areas in the frontal and parietal lobes with a positive relationship to the laboratory measure, as hypothesized. However, it also reveals a negative association spanning broad portions of the fusiform gyrus, as well as vertices showing positive or negative associations scattered across the cortex. The researcher conducting the analysis writes a report detailing the novel findings in the negative direction, emphasizing the stringent methodologies that incorporated two-tailed tests and a strict fdr significance threshold, well below the conventional level $q=0.05$, and concludes (erroneously) that only about 1% of the negative associations are expected to be false positives.
@@
In a study of a novel compound for reducing the severity of relapses in multiple sclerosis (ms), white matter hyperintensities (wmh) are analyzed voxelwise. Subjects are randomly allocated into active compound and placebo groups. Based on prior knowledge and on random population sampling, normality is assumed and a two-sided group comparison is conducted using a two-sample $t$-test. This produces two-tailed p-values calculated from the Student’s $t$-distribution. Control for multiple testing across voxels is achieved using fdr at $q=0.05$. After determining the fdr threshold, the inverse cumulative distribution function of the Student’s $t$-distribution is used to derive the critical threshold $t_{\textsc{fdr}}$. Given the two-tailed nature of the test, symmetrical thresholds $t_{\textsc{fdr}}$ and $t_{\textsc{fdr}}$ are set in an image visualization tool. The results show a notable reduction in the periventricular wmh load in the group that used the active compound, as expected (and hoped), as well as mixed results in various juxtacortical regions, albeit with mostly increases in wmh load in these areas. In the report, the investigator concludes (erroneously) that the drug may amplify the wmh load in juxtacortical areas given that, as fdr was used, no more than 5% of the observed changes in those regions are false positives.
@@
These issues are a consequence of the lack of the subsetting property. Localization errors can be avoided by refraining from making statements or taking actions that could refer to the amount of errors in specific regions when fdr is controlled only globally. The same holds to avoid direction errors. There is, however, a rigorous way to perform localized inferences, and which allows such statements to be made or such actions to be taken. That involves yet another fdr-related procedure, the one proposed by Benjamini2014, presented next.
2.7 Selective inference: bb
The problem of directional inference can be framed as a special case of selective inference on multiple sets of hypotheses. First, consider the $V$ tests partitioned into multiple non-overlapping sets (e.g., all voxels within multiple regions of interest, or all tests partitioned by positive or negative test statistics). A researcher could try to bypass the above issues by applying fdr within each set separately (see Efron2008, for a discussion). In the limiting case in which each set is as small as just one voxel, there is nothing left to be corrected within that set, yet the multiple testing problem would still persist through the existence of multiple sets. The researcher could also try to apply fdr within only select, “promising” sets of tests, such as those that contain small p-values according to some criteria. This, however, would ignore the selection procedure and lead to inflated error rates in those sets that were selected, unless additional adjustments via, e.g., permutations, were used (see Heller2009, for one such procedure).
To address this, Benjamini2014 propose to control neither the error rate within each set, nor the global error rate; instead, the problem is addressed by controlling the average error rate across the sets. To that goal, they develop the following procedure (bb):
  1. 1)
    Compute uncorrected p-values $p_{i}$ for each test.
  2. 2)
    Define the $S$ non-overlapping sets of tests.
  3. 3)
    Apply a screening procedure within each set. Any procedure that depends only on the p-values within that set can be considered, whether controlling fdr or fwer. Procedures such as bh or, equivalently, Simes1986, are fast to compute and are acceptable choices. This is the “first stage” procedure.
  4. 4)
    Count the number $R$ of sets in which at least one test was rejected in the screening procedure.
  5. 5)
    If $R=0$, declare no discoveries and stop. Otherwise, apply an fdr-controlling procedure, such as bh or bky, within each selected set using a modified level $q^{\prime}=qR/S$ to control the expected average fdr over the selected sets. This is the “second stage” procedure.
Benjamini2014 prove that if the selection of a set (in the paper termed family) depends only on the data within that set, then applying an fdr-controlling procedure at level $q^{\prime}=qR/S$ within each selected set guarantees that the average false discovery rate across sets is controlled at $q$. The only requirement is that the screening procedure is local within each set, that is, selection is not a function of p-values from the other sets. Hence, for the Example 1 from the previous section, the problem can be solved by partitioning the statistical map into two disjoint sets, one with the positive test statistics, the other with the negative test statistics, then applying a selection procedure (which can be an fdr procedure) to each side of the map; if nothing survives in either side, the researcher stops. If there is at least one rejection in one ($R=1$) or in each ($R=2$) side, proceed to an fdr procedure in each side, this time using level $q^{\prime}=qR/S$, where $S=2$ given the researcher partitioned the $V$ voxels into two sets. For the Example 2, the problem can be solved by partitioning the map into at least two regions of interest (i.e., two sets, these comprising periventricular and juxtacortical white matter voxels), then if interest is also on making inferences separately across the map, further split each region into positive and negative sides after the test statistics have been computed for each voxel, then subject all such sets of voxels to the bb procedure. In either example, partitioning by the sign of the test statistic works because the sign is independent of the two-tailed p-values; the latter are distributed uniformly between 0 and 1 and, in either side of the map, more extreme test statistics (stronger effects) imply lower p-values. Thus, the requirements for fdr, using either bh or bky, are satisfied. The fdr can then be controlled at the desired level $q$ for the positive and for the negative sides of the map, and further, as in Example 2, for the multiple regions in consideration. Note that, with bb, adjusted p-values can still be computed as described in Section 2.5; only the threshold needs to be modified as $q^{\prime}=qR/S$.